# ONG: Orthogonal Natural Gradient Descent

### v2

@@ 4.5 EKFAC Fisher Approximation
Calculating the exact Fisher information matrix (FIM) and its inverse becomes prohibitively expensive as our model size grows. To make computation tractable, we utilize the Eigenvalue-corrected Kronecker Factorization (EKFAC) George et al. (2018). EKFAC works by providing a diagonal approximation to the FIM not in the raw parameter basis, but in the Kronecker-factored eigenbasis, where a diagonal captures most of the curvature information at a very low cost. Each neural net layer’s Fisher block is approximated as $\boldsymbol{F}=\boldsymbol{A}\otimes\boldsymbol{B}$ with the Kronecker factors $\boldsymbol{A}=\boldsymbol{Q}_{\boldsymbol{A}}\boldsymbol{\Lambda_{A}}\boldsymbol{Q_{A}^{\top}}\text{ and }\boldsymbol{B}=\boldsymbol{Q_{B}}\boldsymbol{\Lambda_{B}}\boldsymbol{Q_{B}^{\top}}$. We track just the diagonal matrices $\boldsymbol{\Lambda_{A}}$ and $\boldsymbol{\Lambda_{B}}$ in these eigenbases, so each mini-batch update just becomes two small eigenbasis transformations plus a diagonal rescaling. We utilize this approximation of the Fisher matrix in our code implementation. We adapt this EKFAC algorithm Vantanapath (2020) in our code implementation for approximating the Fisher matrix, leading to faster and more stable training.
@@ 5.2 Evaluation Criteria
For evaluation purposes, we adapt the same metrics as Bennani et al. (2020): average accuracy and average forgetting. Let $a_{T,k}$ denote the accuracy of the model on the $\mathcal{T}_{k}$ after being trained on the task $\mathcal{T}_{T}$. Average accuracy ($A_{T}$) is the accuracy of the model after its trained on task $\mathcal{T}_{T}$. It is defined as:
@@ 5.4.1 Forgetting and Average Validation Accuracy
Here, we present the average forgetting and average validation accuracy, as described above, of the two baseline methods (OGD and OGD+), as well as our natural gradient extensions to each (ONG and ONG+). Table 1 highlights the average forgetting values and Table 2 highlights the final average validation accuracy.
@@ 5.4.2 Permuted MNIST
Here, we more closely examine our methods’ performance on Permuted MNIST. Table 3 lists the final task-wise test accuracies for each method after the model finishes sequentially training on all tasks. The plots in Figure 1 examine the evolution of validation accuracy during training, specifically how per-task accuracies evolve as the model continues to be trained on newer and newer tasks.
@@ 5.4.3 Rotated MNIST
Here, we more closely examine our methods’ performance on Rotated MNIST. Table 4 lists the final task-wise test accuracies for each method after the model finishes sequentially training on all tasks. The plots in Figure 2 examine the evolution of validation accuracy during training, specifically how per-task accuracies evolve as the model continues to be trained on newer and newer tasks.
@@ 6 Conclusion and Outlook
Our empirical evaluation, however, yielded a counterintuitive yet insightful result: ONG underperforms its Euclidean counterpart on standard continual learning benchmarks. Rather than a simple failure of the method, we interpret this as a key finding of our study: there exists a fundamental tension between Fisher preconditioning and standard orthogonal projections. This discovery suggests that a naive combination of these methods is not only suboptimal but can be detrimental, indicating that the underlying geometries are not immediately compatible. This finding motivates a clear path for future research focused on reconciling these two geometric perspectives. Our primary hypothesis is that the type of projection we utilize must be chosen to be compatible with the descent directions natural gradient descent uses. Immediate future work will therefore involve a deeper analysis of the training dynamics, the evolution of the task-specific gradient subspaces, and the accuracy of the Fisher estimates. Ultimately, the goal is to develop a geometrically-meaningful way of performing "projection" with natural gradients that preserves previously seen gradient directions and enables high accuracy on older tasks.
To that end, incorporating more advanced concepts from Manifold Learning and Information Geometry presents an exciting direction. The concept of parallel transport (see Appendix C for its definition) is particularly promising, as it provides a principled way to move tangent vectors (i.e., gradients) between different points on the parameter manifold Budninskiy et al. (2018); Li and Ma (2022). By using parallel transport to bring gradients from previous tasks into the same tangent space as the current gradient, it may be possible to define a geometrically consistent inner product and basis for projection.
Additional avenues for research include a more rigorous study of the parameterization-invariance property, a primary motivator for this work, to understand which kind of model reparameterizations our method is robust to. Finally, we posit that the limitations we observed may be partially attributable to the high task correlation in benchmarks like Permuted and Rotated MNIST. We hypothesize that the benefits of geometrically-aware methods will become more pronounced in more challenging and realistic continual learning scenarios involving diverse, uncorrelated task sequences. Future work will extend our empirical validation to these more complex domains.
@@ Appendix B Setup
We use the official code implementation of Bennani et al. [2020], which can be obtained here as our starting point for building ONG. We also adapt the EKFAC algorithm provided by Vantanapath [2020] for our efficient Fisher information matrix implementation. Both repositories have a MIT license.

### v3

@@ 4.5 EKFAC Fisher Approximation
Calculating the exact Fisher information matrix (FIM) and its inverse becomes prohibitively expensive as our model size grows. To make computation tractable, we utilize the Eigenvalue-corrected Kronecker Factorization (EKFAC) [George2018EKFAC]. EKFAC works by providing a diagonal approximation to the FIM not in the raw parameter basis, but in the Kronecker-factored eigenbasis, where a diagonal captures most of the curvature information at a very low cost. Each neural net layer’s Fisher block is approximated as $\boldsymbol{F}=\boldsymbol{A}\otimes\boldsymbol{B}$ with the Kronecker factors $\boldsymbol{A}=\boldsymbol{Q}_{\boldsymbol{A}}\boldsymbol{\Lambda_{A}}\boldsymbol{Q_{A}^{\top}}\text{ and }\boldsymbol{B}=\boldsymbol{Q_{B}}\boldsymbol{\Lambda_{B}}\boldsymbol{Q_{B}^{\top}}$. We track just the diagonal matrices $\boldsymbol{\Lambda_{A}}$ and $\boldsymbol{\Lambda_{B}}$ in these eigenbases, so each mini-batch update just becomes two small eigenbasis transformations plus a diagonal rescaling. We utilize this approximation of the Fisher matrix in our code implementation. We adapt this EKFAC algorithm [wiseodd2020naturalgradients] in our code implementation for approximating the Fisher matrix, leading to faster and more stable training.
@@ 5.2 Evaluation Criteria
For evaluation purposes, we adapt the same metrics as [bennani]: average accuracy and average forgetting. Let $a_{T,k}$ denote the accuracy of the model on the $\mathcal{T}_{k}$ after being trained on the task $\mathcal{T}_{T}$. Average accuracy ($A_{T}$) is the accuracy of the model after its trained on task $\mathcal{T}_{T}$. It is defined as:
@@ 5.4.1 Forgetting and Average Validation Accuracy
Here, we present the average forgetting and average validation accuracy, as described above, of the two baseline methods (OGD and OGD+), as well as our natural gradient extensions to each (ONG and ONG+). Table 1 highlights the average forgetting values and Table 2 highlights the final average validation accuracy. As we can see, simply adding a preconditioner to the OGD algorithm leads to extremely high fogetting and lower accuracy. These results suggest that there is some underlying tension between the Euclidian-style projections of OGD with the Fisher-geometry descent directions. While the reason for this discrepancy is unclear, we are currently investigating this phenomena by using crafted datasets and evaluations.
@@ 5.4.2 Permuted MNIST
Here, we more closely examine our methods’ performance on Permuted MNIST. Table 3 lists the final task-wise test accuracies for each method after the model finishes sequentially training on all tasks. The plots in Figure 1 examine the evolution of validation accuracy during training, specifically how per-task accuracies evolve as the model continues to be trained on newer and newer tasks. These finer-grain results show just how quickly the accuracy starts dropping off as we train on newer tasks, and further reinforce the idea that our method of projection is likely misaligned with our metric space of choice.
@@ 5.4.3 Rotated MNIST
Here, we more closely examine our methods’ performance on Rotated MNIST. Table 4 lists the final task-wise test accuracies for each method after the model finishes sequentially training on all tasks. The plots in Figure 2 examine the evolution of validation accuracy during training, specifically how per-task accuracies evolve as the model continues to be trained on newer and newer tasks. While less extreme, we still observe the same sharper dropoff here as well.
@@ 6 Conclusion and Outlook
Our empirical evaluation, however, yielded a counterintuitive yet insightful result: ONG underperforms its Euclidean counterpart on standard continual learning benchmarks. Rather than a simple failure of the method, we interpret this as a key finding of our study: there exists a fundamental tension between Fisher preconditioning and standard orthogonal projections. This discovery suggests that a naive combination of these methods is not only suboptimal but can be detrimental, indicating that the underlying geometry is not immediately compatible. This finding motivates a clear path for future research focused on reconciling these two perspectives. Our primary hypothesis is that the type of projection we utilize must be chosen to be compatible with the descent directions natural gradient descent uses. We are currently investigating on understanding what the exact failure modes are by crafting synthetic, task-datasets and analyzing the training dynamics. Further future work will involve looking more closely at the evolution of the task-specific gradient subspaces, and the accuracy of the Fisher approximation estimates. Ultimately, we aim to utilize these insights and develop a geometrically-meaningful way of performing "projection" with natural gradients that preserves previously seen gradient directions and enables high accuracy on older tasks.
To that end, incorporating more advanced concepts from Manifold Learning and Information Geometry presents an exciting direction. The concept of parallel transport (see Appendix C for its definition) is particularly promising, as it provides a principled way to move tangent vectors (i.e., gradients) between different points on the parameter manifold [budninskiy2018parallel, li2022federated]. By using parallel transport to bring gradients from previous tasks into the same tangent space as the current gradient, it may be possible to define a geometrically consistent inner product and basis for projection.
Additional avenues for research include a more rigorous study of the parameterization-invariance property, a primary motivator for this work, to understand which kind of model reparameterizations our method is robust to. Finally, we posit that the limitations we observed may be partially attributable to the high task correlation in benchmarks like Permuted and Rotated MNIST. We hypothesize that after developing a more roboust, geometrically-aware continual learning method, its benefits will be more pronounced in more challenging and realistic continual learning scenarios involving diverse, uncorrelated task sequences. Future work will extend our empirical validation to these more complex domains.
@@ Appendix B Setup
We use the official code implementation of [bennani], which can be obtained here as our starting point for building ONG. We also adapt the EKFAC algorithm provided by [wiseodd2020naturalgradients] for our efficient Fisher information matrix implementation. Both repositories have a MIT license.